OpenStudy (anonymous):

Choose the procedure you would use to solve the equation 3x − 3 = 35 for x. Add three to both sides, and then multiply both sides by three. Divide both sides by three, and then subtract three from both sides. Multiply both sides by three, and then add three to both sides. Add three to both sides, and then divide both sides by three.

OpenStudy (anonymous):

3x - 3 = 35 3x -3 + 3 = 35 + 3 3x = 38 3x/3 = 38/3 x = 38/3

OpenStudy (anonymous):

so which one will it be then?

OpenStudy (anonymous):

@Deldrickg are u there?

OpenStudy (anonymous):

Anyway, if you look carefully and actually understand what I did above, it is add three to both sides and then divide both sides by 3, so the last option is the answer.